# Counterfactuals As a Means for Evaluating Faithfulness of Attribution Methods in Autoregressive Language Models

### v1

@@ 4.3.2 Gradient $\times$ Input
In gradient $\times$ input (gradinp) method Shrikumar et al. (2016); Denil et al. (2014), we take the dot product of the gradient with the input token embedding $x_{i}$:
@@
By multiplying the gradient with the input embedding, we also account for how much each token is expressed in the attribution score. The Contrastive Gradient $\times$ Input is:
@@ 4.3.3 Erasure
Erasure-based methods measure the importance of each token by erasing it and seeing the effect on the model output (Li et al., 2017). This is achieved by taking the difference of model output with the full input $x$ and part of the input zeroed out, $x_{\neg i}$:
@@ 4.3.4 KernelSHAP
KernelSHAP uses a kernel $\pi_{x}$ that satisfies certain principles when input features are considered agents of a cooperative game in game theory. We use equation 2 in a contrastive way. First we normalize $S_{\phi}^{t}$ and $S_{\phi}^{f}$ by dividing to their $L2$ norm and then subtracting:
@@ 4.3.5 Integrated Gradients
Integrated Gradients (IG) (Sundararajan et al., 2017) is a gradient-based method which addresses the problem of saturation: gradients may get close to zero for a well-fitted function. IG requires a baseline $\mathbf{b}$ as a way of contrasting the given input with information being absent. For input $i$, we compute:
$S^{t}_{IG}=\frac{1}{m}\sum_{k=1}^{m}\nabla_{x_{i}}q\Big{(}y_{t}\Big{|}b+\frac{k}{m}(x-b)\Big{)}\!\cdot\!(x_{i}\!-\!b_{i})\!$ (4)
That is, we average over $m$ gradients, with the inputs to $f_{t}$ being linearly interpolated between the baseline and the original input $x$ in $m$ steps. We then take the dot product of that averaged gradient with the input embedding $\mathbf{x}_{i}$ minus the baseline.
We use zero vector baseline Mudrakarta et al. (2018), and 5 steps. The contrastive case becomes:
@@
Table 2: The mean percentage needed to mask to achieve flipping Gemma-ft’s label or reaching 50 percent masking in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
@@
Table 3: The mean percentage needed to mask to achieve flipping Gemma-it’s label or reaching 50 percent masking in 200 examples of evaluation split in SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
Figure 4: In the top matrix, the average correlation of attribution ranks for the fine-tuned predictor is presented. The middle matrix shows the average correlation of attribution ranks when the predictor is an off-the-shelf model. The bottom matrix displays the difference between the first two, indicating that when using editors, the difference in correlation between the two types of predictors is near zero. However, using other replacement methods results in inconsistent correlations across the two scenarios, likely because these methods produce OOD text for the instruct-tuned model.
@@ 5.1 The Out-of-Distribution problem
Why should we use counterfactuals instead of erasing the important tokens or replacing important tokens with unimportant tokens?
@@
First, we show that our counterfactual generators generate in-distribution text for the predictor. Second, we show that the rankings of faithfulness of attribution methods are consistent with each other when we are using a counterfactual generator for token replacement but these rankings differ when other replacement methods are used.
@@
To achieve our first goal, which is to show that the generated counterfactuals are in-distribution, we use an OOD detection technique to measure what percentage of our generated inputs are OOD. Prominent methods to detect OOD use a threshold and consider anything higher than that threshold as OOD (Chen et al., 2023). To calculate the threshold, for each dataset, we measure the negative log-likelihood (NLL) of the 200 original examples using different predictors—fine-tuned and instruct-tuned—and consider the 99th percentile of the NLLs as the threshold for being OOD. The reason we use NLL to detect OOD is that the type of shift we want to detect is background shift. We can classify the types of OOD data as either semantic or background shift (Arora et al., 2021). Semantic features have a strong correlation with the label, and semantic shift happens when we encounter unseen classes at test time, while background features consist of population-level statistics that do not depend on the label and focus on the style of the text.
In the evaluation of faithfulness by corrupting the input, we are not introducing new labels or classes. We are changing the style of text; therefore, we want to detect background shift. There are two common types of OOD detection methods: calibration and density estimation. Density estimation methods, e.g., perplexity (PPL), outperform calibration methods under background shifts, while the opposite is true under semantic shift. We will use NLL, which is a very similar notion to PPL. An attribution method shows us which tokens are important, and we will replace those tokens in four ways: (i) using an editor to replace the token (our method), (ii) using a token that is considered unimportant semantically (<unk> token and <mask> token), (iii) erasing the tokens, and (iv) zeroing out the attention mask for important tokens without altering the text itself (att-zero). The baselines (ii) through (iv) are similar to previous work (Hase et al., 2021). In Table 1, it is shown that for both fine-tuned and instruct-tuned predictors, the generated counterfactuals are mostly in-distribution. In Table 1, we show the results for the SST-2 dataset and three attribution methods; other attribution methods and datasets are shown in Appendix A. Each number in the table represents the average of five amounts of replacement (10% to 50%) and 200 examples from evaluation sets.
Chen et al. (2023) demonstrate that fine-tuning renders a model insensitive to non-semantic shift. Their research indicates that fine-tuning eliminates pre-trained task-agnostic knowledge about general linguistic properties, which are crucial for detecting non-semantic shift. Our findings align with these observations. When a predictor is fine-tuned for a specific classification task, such as sentiment analysis on the SST-2 dataset, it is optimized to assign the highest probability to the positive token for examples with positive sentiment. Consequently, this fine-tuned model becomes less sensitive to input corruptions. In our experiment, regardless of the replacement method employed, the resulting input tends to remain in-distribution for the fine-tuned predictors. As evidenced in Table 1, under the Gemma-ft columns, the percentage of out-of-distribution (OOD) examples approaches zero. In contrast, Gemma-it, an off-the-shelf model which is not optimized for a specific dataset, exhibits different behavior. When subjected to various input modifications—such as replacing important tokens with semantically neutral ones (e.g., <unk> or <mask> tokens), completely removing tokens, or zeroing out the attention mask for important tokens without altering the text itself—the Gemma-it predictor frequently categorizes these modified inputs as OOD. This disparity in behavior between fine-tuned and off-the-shelf models underscores the impact of task-specific optimization on a model’s sensitivity to input perturbations. Notably, the examples remain in-distribution when modified by the counterfactual generator, even for the instruct-tuned predictor. This observation demonstrates that when we don’t want to change the predictor model and prefer to use an off-the-shelf model as our predictor, using a counterfactual generator is helpful in evaluating the faithfulness of attribution methods.
To achieve our second goal, which is to demonstrate the consistency of the faithfulness rankings of attribution methods regardless of the counterfactual generator used, and the lack of consistency when another replacement method is applied, we use Spearman’s rank correlation, which has been used in previous works (Rong et al., 2022).
We rank the attribution methods’ mask percentage needed to flip the label for each example for all seven replacement methods (our three editors, Erase, <unk>, <mask>, and att-zero), get the correlation of these ranks with each other, and then take the average over 200 examples. We show this for the SST-2 dataset in Figure 4. Other datasets yield similar results and are shown in Appendix B. In the top correlation matrix of Figure 4, these average correlations are shown for the fine-tuned predictor. For the fine-tuned predictor, all replacement methods have high average correlation with each other. The middle matrix in Figure 4 shows these correlations when the predictor models are off-the-shelf instruct-tuned models. For the off-the-shelf predictor, only when a counterfactual generator is used do the ranks have high correlation with each other; other replacement methods have low correlation with counterfactual generators. This is likely because when using an instruct-tuned predictor, replacement methods other than counterfactual generators create OOD inputs. The figure shows the rank of explanations using editors has a higher correlation with unk/erase when the predictor model is fine-tuned. The bottom matrix of 4 is the difference between the first and second matrices. It indicates that the correlation difference between fine-tuned and instruct-tuned predictors is near zero when using editors as the replacement method, but the difference with other corruption methods (unk/erase/mask/att-zero) is significant. This suggests that when evaluating explanations on an off-the-shelf instruct-tuned model, it is crucial not to use corrupted OOD text.
5.2 Analysis of FI methods
@@
In Tables 2 and 3, we show the average masking percent needed (the average percentage of tokens the counterfactual generator should change) to flip the label for fine-tuned and instruct-tuned predictor models, respectively. The masking percent is highly correlated with flip-rate—what percentage of labels each counterfactual generator is able to flip by changing the corrupted tokens. In Appendix C, we show flip-rate for both fine-tuned and instruct-tuned predictor models. Attribution methods that are able to flip the labels with less mask percent (i.e. less change) are also able to flip more labels.
@@
For the fine-tuned predictor (Table 2), gradient norm methods consistently outperform others on the SST-2 and IMDB datasets. In contrast, for AG-News, the Erasure method consistently performs the best or near the best. Our results indicate that straightforward methods, such as gradnorm1, gradnorm2, and Erasure, consistently deliver superior performance regardless of the editor used.
@@
For the instruct-tuned predictor (Table 3), the Erasure method yields the best results for the SST-2 dataset, while gradinp demonstrates the best performance on the IMDB dataset. However, no attribution method consistently outperforms random selection for the AG-News dataset. Overall, these findings suggest that attribution methods are less effective when the model is not fine-tuned for the specific task, indicating the need for cautious application of these methods to pretrained and instruct-tuned language models.
@@ 6 Conclusion
In this work, we designed a faithfulness evaluation protocol based on counterfactual generation. We demonstrated that the efficacy of attribution methods varies between models fine-tuned on our specific dataset and off-the-shelf instruct-tuned models. We showed that counterfactual generators are effective for evaluating feature attribution because they can produce mostly in-distribution text for the predictor model. This approach allows us to separate the evaluation of the model from the evaluation of the attribution method, as the examples used are mostly in-distribution. We also found high consistency between different counterfactual generators and a lack of consistency with other replacement methods, highlighting the importance of being in-distribution, particularly when evaluating attributions on off-the-shelf models. Finally, we used our protocol to compare different attribution methods.
@@ 7 Limitations
Our work is limited in several aspects: First, we rely on generating counterfactuals, for which a strong generative model is needed. Generating counterfactuals, especially for long sequences, is computationally expensive. Second, the counterfactual generator may unintentionally know the artifacts and shortcuts used by the predictor to flip the label, and this could limit the intended application of our approach. Third, we evaluated our protocol for classification datasets; evaluating it for other tasks like translation is left for future work.
@@
Tables 7 and 8 show flip-rates for fine-tuned and instruct-tuned datasets respectively.

### v3

@@ 4.3.2 Gradient $\times$ Input
In gradient $\times$ input (gradinp) method (Shrikumar et al., 2016; Denil et al., 2014), we compute the dot product of the gradient and the input token embedding $x_{i}$:
@@
By multiplying the gradient by the input embedding, we also account for how much each token is expressed in the attribution score. The Contrastive Gradient $\times$ Input is:
@@ 4.3.3 Erasure
Erasure-based methods measure the importance of each token by erasing it and observing the effect on the model output (Li et al., 2017). This is achieved by taking the difference between the model output with the full input $x$ and the model output with the input where token $x_{i}$ is zeroed out, denoted as $x_{\neg i}$:
@@ 4.3.4 KernelSHAP
KernelSHAP uses a kernel $\pi_{x}$ that satisfies certain principles when input features are considered agents of a cooperative game in game theory. We use equation 2 in a contrastive way. First we normalize $S_{\phi}^{t}$ and $S_{\phi}^{f}$ by dividing by their $L2$ norm and then subtracting:
@@ 4.3.5 Integrated Gradients
Integrated Gradients (IG) (Sundararajan et al., 2017) is a gradient-based method which addresses the problem of saturation: gradients may get close to zero for a well-fitted function. IG requires a baseline $\mathbf{b}$ as a way of contrasting the given input with the absence of information. For input $i$, we compute:
$S^{t}_{IG}(x_{i})=\frac{1}{m}\sum_{k=1}^{m}\nabla_{x_{i}}q\Big{(}y_{t}\Big{|}b+\frac{k}{m}(x-b)\Big{)}\!\cdot\!(x_{i}\!-\!b_{i})\!$ (4)
That is, we average over $m$ gradients, with the inputs to $q$ being linearly interpolated between the baseline $b$ and the original input $x$ in $m$ steps. We then take the dot product of that averaged gradient with the input embedding $\mathbf{x}_{i}$ minus the baseline.
We use a zero vector baseline (Mudrakarta et al., 2018) and five steps. The contrastive case becomes:
@@
Table 2: The mean percentage of tokens needed to be masked to achieve flipping Gemma-ft’s label or reaching 50 percent masking in 200 examples from evaluation split of SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
@@
Table 3: The mean percentage of tokens needed to be masked to achieve flipping Gemma-it’s label or reaching 50 percent masking in 200 examples from evaluation split of SST-2, IMDB, and AG-News datasets (lower is better). pythia1, pythia2, and gptj models are used to fill the masks and generate counterfactuals.
Figure 4: The top matrix presents the average correlation of attribution ranks for the fine-tuned predictor. The middle matrix shows the average correlation of attribution ranks when using an off-the-shelf instruct-tuned predictor. The bottom matrix illustrates the difference between the fine-tuned and instruct-tuned models, indicating that when editors are used as the replacement method, the difference in correlation is near zero. In contrast, using other replacement methods (i.e., <unk>, erase, <mask>, att-zero) results in significant inconsistencies between the two predictor types, likely due to the creation of out-of-distribution (OOD) text for the instruct-tuned model.
@@ 5 Results and Discussion
This section begins by addressing the out-of-distribution (OOD) problem, emphasizing the importance of using counterfactual generators over methods like erasing or replacing important tokens with uninformative ones. First, we demonstrate that counterfactual generators produce in-distribution text for both fine-tuned and instruct-tuned models by employing OOD detection techniques focused on background shift, using negative log-likelihood (NLL) as a metric. We find that fine-tuned models are less sensitive to input corruptions, whereas instruct-tuned models classify altered inputs as OOD unless counterfactual generators are used. Second, we show that the faithfulness rankings of attribution methods remain consistent when using counterfactual generators, but differ significantly with other replacement methods, with Spearman’s rank correlation used for evaluation. Lastly, in Subsection 5.2, we analyze feature importance methods by examining the average masking percentage required to flip model predictions. The results indicate that simpler attribution methods perform well for fine-tuned models across different datasets, but are generally less effective for instruct-tuned models, highlighting challenges in applying attribution methods to general-purpose models.
@@ 5.1 The Out-of-Distribution Problem
Why should we use counterfactuals instead of erasing important tokens or replacing them with unimportant ones?
@@
First, we demonstrate that our counterfactual generators produce in-distribution text for the predictor. Second, we show that the faithfulness rankings of attribution methods are consistent when using a counterfactual generator for token replacement, but these rankings differ significantly when other replacement methods are applied.
@@
To achieve our first goal—demonstrating that the generated counterfactuals remain in-distribution—we employ an OOD detection technique to measure the percentage of generated inputs classified as OOD. Prominent OOD detection methods use a threshold, where any input exceeding this threshold is considered OOD (Chen et al., 2023). For each dataset, we compute this threshold by measuring the negative log-likelihood (NLL) of 200 original examples using different predictors (fine-tuned and instruct-tuned) and set the 99th percentile of these NLLs as the OOD threshold. We use NLL for detecting OOD instances because the type of shift we focus on is background shift. OOD data can be categorized as either semantic or background shift (Arora et al., 2021). Semantic features have a strong correlation with the label, and a semantic shift occurs when we encounter unseen classes at test time. In contrast, background shift involves population-level statistics that do not depend on the label and pertain more to the style of the text.
In our evaluation of faithfulness through input corruption, we do not introduce new labels or classes but instead alter the style of the text. Therefore, we focus on detecting background shift. There are two main approaches to OOD detection: calibration-based methods and density estimation methods. Density estimation, such as perplexity (PPL), typically outperforms calibration-based methods when dealing with background shifts, whereas the opposite is true for semantic shifts. Since NLL is closely related to PPL, it is used in our experiments.
An attribution method highlights important tokens, and we replace those tokens using four different approaches: (i) our method of using an editor for token replacement, (ii) replacing tokens with semantically unimportant ones (such as the <unk> or <mask> tokens), (iii) erasing the tokens, and (iv) zeroing out the attention mask for important tokens without altering the text itself (att-zero).
The baseline methods (ii) through (iv) are similar to those used in prior work (Hase et al., 2021). Table 1 shows that, for both fine-tuned and instruct-tuned predictors, the counterfactual generator consistently produces in-distribution text. We present results for the SST-2 dataset using three different attribution methods, while results for other methods and datasets are included in Appendix A. Each value in Table 1 represents the average over five levels of replacement (ranging from 10% to 50%) and 200 examples from evaluation sets.
Chen et al. (2023) show that fine-tuning makes models less sensitive to non-semantic shifts. Fine-tuning reduces task-agnostic, pre-trained knowledge of general linguistic properties that are crucial for detecting such shifts. Our findings align with this observation. When a model is fine-tuned for a specific classification task—such as sentiment analysis on SST-2—it becomes optimized to predict labels with high confidence based on training data. As a result, fine-tuned models are more resilient to input corruptions, regardless of the replacement method. As Table 1 illustrates, under the Gemma-ft columns, the percentage of OOD examples is close to zero for the fine-tuned predictor, regardless of the replacement strategy used.
In contrast, Gemma-it, an off-the-shelf model not fine-tuned for a specific dataset, behaves differently. When important tokens are replaced with semantically neutral tokens (e.g., <unk> or <mask>), erased, or subjected to attention masking without textual alteration, the Gemma-it predictor frequently categorizes these modified inputs as OOD. This difference in behavior between fine-tuned and off-the-shelf models highlights the impact of task-specific optimization on a model’s sensitivity to input perturbations. However, when using the counterfactual generator to modify inputs, the modified examples remain in-distribution even for the instruct-tuned predictor. This demonstrates that counterfactual generators are useful for evaluating the faithfulness of attribution methods, especially when using off-the-shelf models where modification without creating OOD inputs is essential.
To achieve our second goal—demonstrating that faithfulness rankings of attribution methods remain consistent when using a counterfactual generator, and inconsistent with other replacement methods—we use Spearman’s rank correlation, as in prior studies (Rong et al., 2022). For each example, we rank attribution methods based on the percentage of the mask required to flip the label. We then compute the correlation between these rankings across seven replacement methods (three editors, Erase, <unk>, <mask>, and att-zero), averaging the results over 200 examples.
Figure 4 presents this analysis for the SST-2 dataset. Similar results for other datasets are shown in Appendix B. The top matrix in Figure 4 shows average correlation results for the fine-tuned predictor, where all replacement methods exhibit high average correlations with one another. The middle matrix depicts the correlations for the off-the-shelf instruct-tuned model. In this case, only when using a counterfactual generator do the rankings show high correlation, while other replacement methods produce low correlations. This is likely due to OOD inputs being created by non-counterfactual replacement methods.
The bottom matrix in Figure 4 illustrates the difference between the first and second matrices. It shows that the correlation difference between fine-tuned and instruct-tuned predictors is negligible when editors are used for token replacement. However, significant differences arise when other methods (e.g., <unk>/Erase/<mask>/att-zero) are employed, underscoring the importance of avoiding OOD inputs when evaluating explanations for off-the-shelf models.
5.2 Analysis of Feature Importance Methods
@@
Tables 2 and 3 present the average masking percentage required (i.e., the average percentage of tokens that the counterfactual generator needs to modify) to flip the prediction for fine-tuned and instruct-tuned predictor models, respectively. This masking percentage is strongly correlated with the flip rate—the proportion of labels the counterfactual generator successfully flips by altering the identified important tokens. In Appendix C, we provide a detailed analysis of the flip rate for both fine-tuned and instruct-tuned models. Attribution methods that require less masking to flip a label tend to also have a higher flip rate, indicating greater faithfulness.
@@
For the fine-tuned predictor (Table 2), gradient norm-based methods consistently outperform others on the SST-2 and IMDB datasets. However, for AG-News, the Erasure method consistently performs the best or near the best. These results suggest that simpler attribution methods, such as gradnorm1, gradnorm2, and Erasure, provide superior performance across different editors and datasets.
@@
For the instruct-tuned predictor (Table 3), the Erasure method achieves the best results for the SST-2 dataset, while the gradinp method performs best on the IMDB dataset. Interestingly, for the AG-News dataset, no attribution method consistently outperforms random selection, highlighting the challenges of applying attribution methods to general-purpose models. These findings emphasize that attribution methods are generally less effective when the model is not fine-tuned for a specific task, suggesting the need for careful consideration when applying these methods to pretrained or instruct-tuned models.
@@ 6 Conclusion
In this work, we introduced a faithfulness evaluation protocol based on counterfactual generation to assess the performance of attribution methods. We demonstrated that the efficacy of these methods differs significantly between models that are fine-tuned on a specific dataset and those that are instruct-tuned and used off-the-shelf. Our results showed that counterfactual generators are highly effective for evaluating feature attribution, as they consistently produce in-distribution text for the predictor model. This ensures a clearer separation between the evaluation of the model itself and the attribution method, with minimal influence from out-of-distribution examples.
Moreover, our findings revealed strong consistency across different counterfactual generators, while other token replacement methods lacked this consistency. This underscores the importance of maintaining in-distribution inputs, especially when evaluating attribution methods for off-the-shelf models. Lastly, we applied our protocol to compare various attribution methods, providing insights into their relative effectiveness across different model configurations.
@@ 7 Limitations
Our work has several limitations. First, it relies on the generation of counterfactuals, which necessitates the use of a robust generative model. This process, particularly for long sequences, can be computationally expensive and may limit the scalability of our approach. Second, there is a risk that the counterfactual generator may unintentionally exploit artifacts or shortcuts used by the predictor model to flip the label, which could undermine the broader applicability of our method. Finally, our protocol has been tested exclusively on classification tasks; extending it to other tasks, such as machine translation or question answering, remains an area for future exploration.
@@
Tables 7 and 8 show flip-rates for fine-tuned and instruct-tuned predictor models respectively.